# Do PAC-Learners Learn the Marginal Distribution?

### v2

@@ 2.5 Further Related Work
We first describe the construction, which is similar in spirit to the UBME counterexample of DKRZ [11]. Let $X=\{0,1,2\}^{\mathbb{N}}$, and $\{\epsilon_{i}\}_{i\in\mathbb{N}}$ be the sequence $\varepsilon_{i}\coloneqq\frac{1}{\log(i+255)}$. For each $i\in\mathbb{N}$, let $f_{i}$ denote the indicator that $x_{i}=0$:
@@
$\Pr[x_{j}=0]=\Pr[x_{j}=1]=\epsilon_{j},\leavevmode\nobreak\ \text{and }\Pr[x_{j}=2]=1-2\epsilon_{j}.$
@@
Fix any $D_{i}\in\mathscr{P}$ and $f_{j}\in H$. First, observe that by a Chernoff bound, $f_{j}\in G_{T,D_{i}}$ with probability at least $1-\delta$. Second, notice that since $\mu_{i}\coloneqq\mathbb{E}_{D_{i}}[f_{i}]=1/2$ and any other $\mu_{j}\leq 1/8$, for all functions $f_{\ell}\in G_{T,D_{i}}$, we are assured that $\mu_{\ell}^{T}<1/3$ if and only if $\ell\neq i$. Since this condition exactly determines the distances between hypotheses, $\tilde{d}_{T}$ has zero error across $G_{T,D_{i}}$ as desired. ∎
@@
by construction. Moreover, by our choice of $\{\epsilon_{j}\}_{j\in\mathbb{N}}$, we have $k\leq O(2^{1/\varepsilon})$ which gives the desired bound on metric entropy. ∎
@@
Since we chose $\epsilon_{j}=\frac{1}{\log(j+255)}$ to decay sufficiently slowly, the sum in the exponent diverges. Thus for any fixed $t$ there is a sufficiently large $n$ for which $\Pr[E_{W}:|W|=1]\leq 1/2$ which completes the proof. ∎
5 Density Estimation, Uniform Convergence, and UBME
@@
samples, where $\varepsilon^{\prime}=O(\varepsilon)$, $\delta^{\prime}=O(\frac{\delta}{\Pi_{H}(\varepsilon/4,1/2)})$.
@@
where $D$ is the adversary’s choice of distribution. We call $(C,c)$ a covering-map pair. This type of structure is useful since it allows us to apply Yatracos’ trick [13] of extending empirical estimates on $C$ (which is finite), to good estimates on all of $H$ via the covering map. With this in mind we prove the backward direction of Theorem 5.1 which follows similarly to Theorem 4.3, replacing the non-uniform cover with a true covering-map pair.
Proof of Theorem 5.1: DE+UBME $\implies$ UE.
@@
So long as $(C_{D^{\prime}},c_{D^{\prime}})$ is truly an $(\varepsilon/2)$-covering-map pair under $D$, the error of $h$ and $c(h)$ can differ by at most $\varepsilon/2$. Thus if the empirical errors on $C_{D^{\prime}}$ are accurate up to $\varepsilon/2$, $\mathcal{E}_{S}$ is within $\varepsilon$ of the true risk on all hypotheses by construction. By a Chernoff and Union bound this latter guarantee also holds with probability at least $21-\delta/$ 2, which completes the result. ∎
To prove the forward direction of Theorem 5.1, we will need the following related lemma of HKLM [18] that shows how to build a finite cover of any learnable class.
Lemma 5.2 (Learnable Classes are Coverable [18, Theorem G.9]).
@@
Note that this algorithm does not produce a corresponding covering map (indeed this would imply necessity of density estimation, which we refuted in Section 3). In this case, we will use the uniform estimator to build a corresponding covering map, from which it is then easy to deduce a density estimate.
Proof of Theorem 5.1: UE $\implies$ DE+UBME.
@@
Such a distribution is guaranteed to exist by validity of the estimates, and satisfies $TV_{H\Delta H}(D,D^{\prime})\leq\varepsilon$ by construction. Union bounding over all events, the algorithm succeeds with probability at least $1-\delta$ which completes the proof. ∎
@@
Proof.
If $(X,H)$ is PAC-learnable, then it satisfies uniform convergence [2, 3] which implies the existence of a density estimator by Theorem 5.1. If $(X,H)$ has a density estimator and has UBME, then it is PAC-learnable by Proposition 3.1. ∎
We remark that in the distribution-free case, one does not have to factor through uniform estimation, and could instead use a more classical approach by arguing $H\Delta H$ has finite VC and therefore satisfies fast convergence of empirical densities in $H\Delta H$-distance (see [48, Theorem 4.9]). This approach fails in our setting since $H\Delta H$ need not have finite VC.
@@ 5.2 Uniform Convergence and UBME
Proof.
Containment is immediate from the characterization of DE + UBME as uniform estimation (Theorem 5.1). Strictness follows from Benedek and Itai’s [9] example of a learnable class that fails uniform convergence. Let $D$ be the uniform distribution over $[0,1]$, and let $H$ consist of all sets of finite support and the all $1$’s function. It is easy to see uniform convergence fails, since choosing $h=\vec{1}$, for any finite sample $T$ there is always a completely consistent hypothesis with error $1$, namely the indicator $\mathbf{1}_{T}$. On the other hand, since $\mathscr{P}$ is just a single distribution, density estimation is trivial. ∎
@@
as desired. Since any fixed $h\in H$ lies in $G_{T,D}$ with probability at least $1-\delta$ by a Chernoff bound and distances in $C$ are also well estimated with probability at least $1-\delta$, this gives the desired estimator. ∎
Note that if one is not interested in relative sample complexities, this subsumes Proposition 4.2 since any PAC-learnable class has UBME [9]. Proposition 5.5 also implies that the counter-example of DKRZ [11] shows weak estimation fails sufficiency as well. However, our construction in the previous section gives a stronger separation since the above approach is inherently lossy and cannot handle $\varepsilon=0$.
Acknowledgements
We thank Shay Moran for many interesting discussions surrounding the topic of this work and anonymous reviewers for pointing out connections with density estimation, and suggesting content rephrasing and improved naming conventions.

### v3

@@ Acknowledgements
We first describe the construction, which is similar in spirit to the UBME counterexample of Dudley et al. (1994). Let $X=\{0,1,2\}^{\mathbb{N}}$, and $\{\epsilon_{i}\}_{i\in\mathbb{N}}$ be the sequence $\varepsilon_{i}\coloneqq\frac{1}{\log(i+255)}$. For each $i\in\mathbb{N}$, let $f_{i}$ denote the indicator that $x_{i}=0$:
@@
$\Pr[x_{j}=0]=\Pr[x_{j}=1]=\epsilon_{j},~{}\text{and }\Pr[x_{j}=2]=1-2\epsilon_{j}.$
@@
Fix any $D_{i}\in\mathscr{P}$ and $f_{j}\in H$. First, observe that by a Chernoff bound, $f_{j}\in G_{T,D_{i}}$ with probability at least $1-\delta$. Second, notice that since $\mu_{i}\coloneqq\mathbb{E}_{D_{i}}[f_{i}]=1/2$ and any other $\mu_{j}\leq 1/8$, for all functions $f_{\ell}\in G_{T,D_{i}}$, we are assured that $\mu_{\ell}^{T}<1/3$ if and only if $\ell\neq i$. Since this condition exactly determines the distances between hypotheses, $\tilde{d}_{T}$ has zero error across $G_{T,D_{i}}$ as desired.
@@
by construction. Moreover, by our choice of $\{\epsilon_{j}\}_{j\in\mathbb{N}}$, we have $k\leq O(2^{1/\varepsilon})$ which gives the desired bound on metric entropy.
@@
Since we chose $\epsilon_{j}=\frac{1}{\log(j+255)}$ to decay sufficiently slowly, the sum in the exponent diverges. Thus for any fixed $t$ there is a sufficiently large $n$ for which $\Pr[E_{W}:|W|=1]\leq 1/2$ which completes the proof.
[…]
@@
samples, where $\varepsilon^{\prime}=O(\varepsilon)$, $\delta^{\prime}=O(\frac{\delta}{\Pi_{H}(n_{UE}(\varepsilon/4,\delta/2))})$.
@@
where $D$ is the adversary’s choice of distribution. We call $(C,c)$ a covering-map pair. This type of structure is useful since it allows us to apply Yatracos (1985)’s trick of extending empirical estimates on $C$ (which is finite), to good estimates on all of $H$ via the covering map. With this in mind we prove the backward direction of Theorem 6.1 which follows similarly to Theorem 5.3, replacing the non-uniform cover with a true covering-map pair.
Proof 6.2 (Proof of Theorem 6.1: DE+UBME $\implies$ UE).
@@
So long as $(C_{D^{\prime}},c_{D^{\prime}})$ is truly an $(\varepsilon/2)$-covering-map pair under $D$, the error of $h$ and $c(h)$ can differ by at most $\varepsilon/2$. Thus if the empirical errors on $C_{D^{\prime}}$ are accurate up to $\varepsilon/2$, $\mathcal{E}_{S}$ is within $\varepsilon$ of the true risk on all hypotheses by construction. By a Chernoff and Union bound this latter guarantee also holds with probability at least $21-\delta/$ 2, which completes the result.
To prove the forward direction of Theorem 6.1, we will need the following related lemma of Hopkins et al. (2024) that shows how to build a finite cover of any learnable class.
Lemma 6.3 (Learnable Classes are Coverable (Hopkins et al., 2024, Theorem G.9)).
@@
Note that this algorithm does not produce a corresponding covering map (indeed this would imply necessity of density estimation, which we refuted in Section 3.1). In this case, we will use the uniform estimator to build a corresponding covering map, from which it is then easy to deduce a density estimate.
Proof 6.4 (Proof of Theorem 6.1: UE $\implies$ DE+UBME).
@@
Such a distribution is guaranteed to exist by validity of the estimates, and satisfies $TV_{H\Delta H}(D,D^{\prime})\leq\varepsilon$ by construction. Union bounding over all events, the algorithm succeeds with probability at least $1-\delta$ which completes the proof.
@@
Proof 6.6.
If $(X,H)$ is PAC-learnable, then it satisfies uniform convergence (Vapnik and Chervonenkis, 1974; Blumer et al., 1989) which implies the existence of a density estimator by Theorem 6.1. If $(X,H)$ has a density estimator and has UBME, then it is PAC-learnable by Proposition 4.1.
We remark that in the distribution-free case, one does not have to factor through uniform estimation, and could instead use a more classical approach by arguing $H\Delta H$ has finite VC and therefore satisfies fast convergence of empirical densities in $H\Delta H$-distance (see e.g. (Anthony et al., 1999, Theorem 4.9)). This approach fails in our setting since $H\Delta H$ need not have finite VC.
@@ 6.2 Uniform Convergence and UBME
Proof 6.8.
Containment is immediate from the characterization of DE + UBME as uniform estimation (Theorem 6.1). Strictness follows from Benedek and Itai (1991)’s example of a learnable class that fails uniform convergence. Let $D$ be the uniform distribution over $[0,1]$, and let $H$ consist of all sets of finite support and the all $1$’s function. It is easy to see uniform convergence fails, since choosing $h=\vec{1}$, for any finite sample $T$ there is always a completely consistent hypothesis with error $1$, namely the indicator $\mathbf{1}_{T}$. On the other hand, since $\mathscr{P}$ is just a single distribution, density estimation is trivial.
@@
as desired. Since any fixed $h\in H$ lies in $G_{T,D}$ with probability at least $1-\delta$ by a Chernoff bound and distances in $C$ are also well estimated with probability at least $1-\delta$, this gives the desired estimator.
Note that if one is not interested in relative sample complexities, this subsumes Proposition 5.1 since any PAC-learnable class has UBME (Benedek and Itai, 1991). Proposition 6.9 also implies that the counter-example of Dudley et al. (1994) shows weak estimation fails sufficiency as well. However, our construction in the previous section gives a stronger separation since the above approach is inherently lossy and cannot handle $\varepsilon=0$.
[…]
We thank Shay Moran for many interesting discussions surrounding the topic of this work and anonymous reviewers for pointing out connections with density estimation and suggesting improved naming conventions.